# Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems

### v1

@@ Abstract
A quantum kinetic theory of the linear response to an electric field is provided from a controlled expansion of the Keldysh theory at leading order, for a general multiband electron system with weak scalar disorder. The response is uniquely partitioned into intraband distribution functions and interband quantum coherences. A new explicit formula is provided for the latter, outlining their true nature as local slave quantities, for which the mesoscopic gradients of the former is uncovered as a new source term. The precise connection with the Kubo formula with vertex corrections in the ladder approximation is established. A pedagogical application to spin-orbit torque theory in the two dimensional electron gas demonstrates the striking efficiency of the new formalism. Some important implications for the theory of orbital transport are also briefly discussed.
@@
Introduction– Linear response theory based on the Kubo formula[1, 2], usually expressed in more practical forms of the Kubo-Bastin or Kubo-Streda types [3, 4, 5, 6, 7], is instrumental to our understanding of the electron Fermi liquid properties in solids. Its application to the theory of quantum Hall effect [8, 9] provides a salient illustration of its power in this context. Berry phase effects [10] are fully included, as seen for instance in the expressions one can derive for the intrinsic anomalous Hall effect (AHE) [11], spin Hall effect (SHE) [12] or orbital Hall effect (0HE) [13, 14] conductivities. It is also naturally amenable to diagrammatic expansions [15, 16, 17], allowing to account for disorder in a controlled manner. However, concrete Kubo based calculations often incorporate disorder effects only through a constant broadening approximation (CBA) [18, 19]. Such an uncontrolled approximation, completely neglecting vertex corrections, is often problematic [20, 21]. The Kubo formula also enables direct computations from first principles, see for instance [22, 23, 24, 25], leading to quantitative predictions for actual material systems. However, in many systems under current intense investigation, such as spintronic [26], orbitronic [27] or topological [28] nanostructured materials and devices, one is primarily interested in disorder averaged responses in finite mesoscopic samples. Full quantum simulations in such cases, specially if one wants to go beyond the CBA, become quite involved, even with the emergence of efficient real space numerical formulations [29, 30, 31, 32, 33]. Alternatively, linear response can be derived from semiclassical approaches [34, 35, 36, 37], which can be seen as some attempts to incorporate leading quantum geometric effects and disorder corrections to the simple quasiparticle transport picture provided by the Boltzmann equation (BE) [38]. While this is in principle more suited to model spatially inhomogeneous systems, these earlier works did not provide a full account of the effects of gradients on mesoscopic length scales, while incorporating on an equal footing leading Berry phase effects and vertex corrections associated with weak disorder.
@@
In this letter, we provide a consistent quantum kinetic theory of the linear response to a nearly constant and uniform electric field ${\bm{E}}$, for a multiband electron system modeled by a low energy Hamiltonian, in presence of weak scalar disorder. This has been obtained from a controlled semiclassical expansion of the Schwinger-Keldysh theory [39, 40, 41], in a gauge invariant formulation in phase space [42] now extended beyond the clean limit. With the details of all the derivations to be found here [43], we first expose our general results in terms of intraband and interband responses, uncovering mesoscopic gradients of the former as a previously overlooked source of the latter. With the help of the Matsubara formalism [15, 16], we then establish the equivalence of our results with the Kubo formula with vertex correction in the ladder approximation, for the spatially homogeneous case. As a pedagogical example, we consider anew the theory of spin-orbit torque (SOT) in the two-dimensional electron gas (2DEG), and uncover a new instance of intrinsic response as a Fermi surface property, reminiscent of the classical result obtained by Haldane in the case of the AHE [44]. In conclusion, we briefly discuss future extensions and we outline some implications for the theory of the recently observed orbital edge accumulation [45, 46] and orbital torque [47, & references therein].
@@
General formalism – Our main new result is a general expression for the interband part of the density matrix i.e., $\hat{\rho}_{inter}^{(1)}=\sum_{n\neq m}\hat{\rho}_{nm}^{(1)}$, in linear response to ${\bm{E}}$, with
$\hat{\rho}_{nm}^{(1)}$ $=-\hbar\frac{f_{n}^{(0)}-f_{m}^{(0)}}{\varepsilon_{n}-\varepsilon_{m}}e{\bm{E}}\cdot\hat{\mathcal{\bm{A}}}_{nm}-\hbar\partial_{\bm{x}}\frac{f_{n}^{(1)}+f_{m}^{(1)}}{2}\cdot\hat{\mathcal{\bm{A}}}_{nm}$
@@
×$\!\!\!\!\!\!\!\!\!\!\!\cdots\times\left[\delta\left(\underaccent{\tilde}{\varepsilon}_{q}\!-\!\varepsilon_{n}\right)\left(\underaccent{\tilde}{f}_{q}^{(1)}\!\!-\!\!f_{n}^{(1)}\right)+\delta\left(\underaccent{\tilde}{\varepsilon}_{q}\!-\!\varepsilon_{m}\right)\left(\underaccent{\tilde}{f}_{q}^{(1)}\!\!-\!\!f_{m}^{(1)}\right)\right],$
while the intraband part $\hat{\rho}_{intra}^{(1)}=\sum_{n}f^{(1)}_{n}\hat{P}_{n}$ is controlled by a system of linear BEs
$\partial_{t}f_{n}^{(1)}+\partial_{\bm{p}}\varepsilon_{n}\cdot\partial_{\bm{x}}f_{n}^{(1)}+(e{\bm{E}})\cdot\partial_{\bm{p}}\varepsilon_{n}\partial_{\varepsilon}f_{n}^{(0)}=C_{n}[{\bm{f}}^{(1)}].$ (2)
@@
To be more precise, $\hat{\rho}^{(1)}({\bm{x}},t;{\bm{p}})\equiv\hat{\rho}^{(1)}=\hat{\rho}^{(1)}_{intra}+\hat{\rho}^{(1)}_{inter}$ as determined by Eqs.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2) is the gauge invariant Wigner function i.e., the gauge invariant Wigner transform of the density matrix [48]. Once $\hat{\rho}^{(1)}$ is obtained, the expectation value of any single particle observable $\smash{{\hat{O}}}$, in the form of its local density $\smash{o^{(1)}({\bm{x}},t)}$, is obtained from its Weyl symbol $\smash{{\hat{o}}({\bm{x}},{\bm{p}})}$ as $o^{(1)}({\bm{x}},t)=\int\frac{d^{d}p}{(2\pi\hbar)^{d}}\ \mathrm{Tr}[{\hat{\rho}}^{(1)}({\bm{x}},t;{\bm{p}})\ {\hat{o}}({\bm{x}},{\bm{p}})]$. Equations (Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2) determine the linear response of a non interacting system of electrons with charge $e(<0)$, specified by a multiband low energy Hamiltonian matrix i.e., ${\hat{h}}={\hat{h}}({\bm{p}})$, with ${\bm{p}}$ the kinetic crystal pseudo momentum. This matrix is acting on the Hilbert space of dimension $N$ of the internal degrees of freedom (spin, orbital), and it may equally correspond to a “toy” model or to a sophisticated Wannier tight-binding Hamiltonian [49, 50, 51]. It has in general $M\leq N$ distinct eigenvalues, the band energies, introduced as $\varepsilon_{n}\equiv\varepsilon_{n}({\bm{p}})$. The projectors into the corresponding eigenspaces, introduced as $\hat{P}_{n}\equiv\hat{P}_{n}({\bm{p}})$, read [52] $\hat{P}_{n}=\prod_{m}\frac{\hat{h}-\varepsilon_{m}\hat{\gamma}_{0}}{\varepsilon_{n}-\varepsilon_{m}}$, with $\hat{\gamma}_{0}$ the identity matrix. We further define $\hat{\mathcal{\bm{A}}}_{nm}={\bf i}\hat{P}_{n}\partial_{\bm{p}}\hat{P}_{m}$, which appears in the first line of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). In the special case of two non degenerate bands, it is immediate to verify that we have $(\hat{\mathcal{\bm{A}}}_{nm})^{i}={\mathcal{A}}^{i}_{nm}|n\rangle\langle m|$, with ${\mathcal{A}}^{i}_{nm}={\bf i}\langle n|\partial_{p^{i}}m\rangle$ the previously introduced abelian interband Berry connection [53]. Hence, we recognize $\hat{\mathcal{\bm{A}}}_{nm}$ as a generalized non abelian interband Berry connection, defined in a manifestly $SU(N)$ gauge invariant way. Also appearing in Eqs.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2), are the linear changes in the intraband distribution functions i.e., $f_{n}^{(1)}\equiv f_{n}^{(1)}({\bm{x}},t;{\bm{p}})$, reducing in equilibrium to the Fermi-Dirac distribution function $f_{n}^{(0)}\equiv n_{\scriptscriptstyle FD}(\varepsilon_{n})$. In the present work, we purposely limit our consideration of disorder to the most simple and generic model i.e., a dilute random distribution of scalar point scattering potentials treated at the level of the self-consistent second Born approximation after configuration averaging. It is parameterized by the density $n_{i}$ of defects and by a momentum independent scattering amplitude $v_{0}$. The collision integral appearing on the right-hand-side of the BE (2) reads
@@
$\!\!\!\!C_{n}[{\bm{f}}^{(1)}]\!=\!\!\sum_{m}\!\!\int\!\!\!\frac{d^{d}\underaccent{\tilde}{p}}{(2\pi\hbar)^{d}}\delta\left(\underaccent{\tilde}{\varepsilon}_{m}-\varepsilon_{n}\right)W_{nm}\left(\underaccent{\tilde}{f}_{m}^{(1)}-f_{n}^{(1)}\right)\!,$ (3)
@@
$W_{nm}\equiv W_{nm}({\bm{p}},\underaccent{\tilde}{\bm p})=\frac{2\pi n_{i}v_{0}^{2}}{\hbar}{\rm Tr}\left[\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{m}\right].$ (4)
@@
It has to be noted that we systematically use a tilde under-accent to indicate that a certain quantity depends on the integration momentum variable $\underaccent{\tilde}{\bm p}$. After introducing $\tau\sim\hbar/(n_{i}v_{0}^{2})$ as the momentum relaxation time scale, we see that $\xi_{i}=1/\tau$ defines a natural scaling parameter for the disorder, while $\hbar$ is of course the formal expansion parameter for the quantum corrections to the classical kinetics.
@@
While the commonly considered decomposition of the Kubo-Streda type formulas into Fermi sea and Fermi surface terms [5, 7] is not physically significant in itself [44], the quantum kinetic approach leads naturally to a unique and gauge invariant split at the Wigner function level i.e., $\hat{\rho}^{(1)}=\hat{\rho}^{(1)}_{intra}+\hat{\rho}^{(1)}_{inter}$. Its physical meaning is further illuminated by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), establishing the interband response components as slave variables locally and uniquely determined by the quantum geometry and the solutions of the BE (2). Hence, our formalism establishes that direct interband contributions to observable densities are not quantities susceptible to advective or diffusive transport in real space. Further looking into the different terms on the right-hand side (RHS) of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we recognize the first term as of purely quantum geometric origin and of order $\hbar\xi_{i}^{0}$, which will yield for instance the intrinsic contribution to the AHE [54, 11] which can be expressed in terms of the Berry curvature. If we look now at the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), since $f_{n}^{(1)}\sim\xi_{i}^{-1}$, as it can be immediately verified by direct inspection of Eqs.(2-4), we see that it is also of order $\hbar\xi_{i}^{0}$ due to the compensating factor $(n_{i}v_{0}^{2})/\hbar\sim\xi_{i}$ in front of the momentum integral. Hence, even if this contribution is disorder mediated, it is ultimately independent of its strength. As demonstrated in the next section, this corresponds to the vertex correction in the ladder approximation to the interband part of the Kubo formula, for which we provide here a new explicit expression at the Wigner function level, in terms of the solutions to the BE (2). As previously mentioned, even recent theoretical studies rely most often on a simplified version of the Kubo-Streda formula at the CBA level, with the leading disorder correction to the interband response found to be of order $\hbar\xi_{i}$ [55], which is qualitatively wrong and totally misses the true leading effect of disorder. As for the second term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), it indicates that mesoscopic gradients in spatially inhomogeneous solutions of the BE (2) are also a source of interband quantum coherences, for which we provide here an explicit expression. To the best of our knowledge, this is an entirely new finding. Such gradients typically appear near boundaries and interfaces on the length scale of the electron mean-free-path [56, 57, 58], hence $\partial_{\bm{x}}f_{n}^{(1)}\sim\xi_{i}^{0}$, which demonstrates that gradient induced coherences are of the same leading order as the two other contributions to the interband response. As a final remark, regarding now the intraband response, it shall be noted that there is no Berry curvature corrections to the BE (2), at the leading order $f_{n}^{(1)}\sim\xi_{i}^{-1}$. There are however sub-leading corrections to the intraband linear response at order $\xi_{i}^{0}$ which involve the quantum geometry, and are related to the intrinsic side-jump and skew-scattering contributions found with the semiclassical approach [35, 36, 37]. This is beyond the scope of the present letter, since a consistent account requires to go beyond the ladder approximation [59], and will be provided within our general framework in a future work.
@@
Correspondence with the Kubo formula– It is well established that the intraband part of the linear response controlled by the BE (2), in the constant and homogeneous case, is identical to the intraband part of the Kubo formula with the vertex correction in the ladder approximation [15, 60]. In this section, we establish a similar equivalence for the interband response as given by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). According to the Kubo formula at finite temperature in the constant and translationally invariant limits [15, 16], the expectation value of an observable ${\hat{o}}(\bm{p})\equiv{\hat{o}}$ reads $o^{(1)}=-eE^{i}\int_{p}\lim_{\omega\rightarrow 0}R^{i}(\omega)/({\bf i}\omega)$, with $\int_{p}\equiv\int\frac{{\rm d}^{d}{{p}}}{(2\pi\hbar)^{d}}$ and with $R^{i}({\bm{p}},\omega)\equiv R^{i}(\omega)$ being the analytical continuation (AC) of the Matsubara response function i.e., $R^{i}(\omega)={\cal R}^{i}(\omega_{\nu}\rightarrow-{\bf i}(\hbar\omega+{\bf i}0^{+}))$, with
@@
with $^{\hat{j}}^{i}=\partial_{p^{i}}\hat{h}$ the bare current vertex. From the corresponding expression of $o^{(1)}$ in terms of the density matrix, we immediately establish the correspondance
$\!\!{\hat{\rho}}^{(1)}_{nm}=-eE^{i}\!\lim_{\omega\rightarrow 0}\frac{\left[\sum_{\epsilon_{\mu}}\!\!\hat{P}_{n}\hat{\mathcal{G}}(\epsilon_{\mu}+\omega_{\nu}){\hat{J}}^{i}\hat{\mathcal{G}}(\epsilon_{\mu})\hat{P}_{m}\right]_{\scriptscriptstyle AC}}{{\bf i}\omega},$ (7)
@@
We can also write ${\hat{J}}^{i}=\sum_{n}J^{i}_{n}\hat{P}_{n}+\sum_{n\neq m}\hat{P}_{n}{\hat{J}}^{i}_{nm}\hat{P}_{m}$. At leading order in $\xi_{i}$, we then obtain the band projected part of Eq.(6) as

### v2

@@ Abstract
A quantum kinetic theory of the linear response to an electric field is provided from a controlled expansion of the Keldysh theory at leading order, for a multiband electron system with weak scalar disorder. The response is uniquely partitioned into intraband distribution functions and interband quantum coherences. A new explicit formula is provided for the latter, outlining their true nature as local slave quantities, for which the mesoscopic gradients of the former is uncovered as a source term, opening a new research area. The precise connection with the Kubo formula with vertex corrections in the ladder approximation is established. A pedagogical application to spin-orbit torque theory in the two dimensional electron gas demonstrates the striking efficiency of the new formalism. Some important implications for the theory of orbital transport are also briefly discussed.
@@
Introduction– Linear response theory based on the Kubo formula[1, 2], usually expressed in more practical forms of the Kubo-Bastin or Kubo-Streda types [3, 4, 5, 6, 7], is instrumental to our understanding of the electron Fermi liquid properties in solids. Its application to the theory of the quantum Hall effect [8, 9] provides a salient illustration of its power in this context. Berry phase effects [10] are fully included, for instance in the expressions one can derive for the intrinsic anomalous Hall effect (AHE) [11], spin Hall effect (SHE) [12] or orbital Hall effect (OHE) [13, 14] conductivities. It is also naturally amenable to diagrammatic expansions [15, 16, 17], allowing to account for disorder in a controlled manner. However, concrete Kubo based calculations often incorporate disorder effects only through a constant broadening approximation (CBA) [18, 19]. Such an uncontrolled approximation, completely neglecting vertex corrections, is often problematic [20, 21]. The Kubo formula also enables direct computations from first principles, see for instance [22, 23, 24, 25], leading to quantitative predictions for actual material systems. However, in many systems under current intense investigation, such as spintronic [26], orbitronic [27] or topological [28] nanostructured materials and devices, one is primarily interested in disorder averaged responses in finite mesoscopic samples. Full quantum simulations in such cases, specially if one wants to go beyond the CBA, become quite involved, even with the emergence of efficient real space numerical formulations [29, 30, 31, 32, 33]. Alternatively, the linear response can be derived from quasiclassical [34], semiclassical [35, 36, 37] or density matrix equation of motion [38, 39] approaches, which can be seen as attempts to incorporate leading quantum geometric and disorder corrections to the simple quasiparticle transport picture of the Boltzmann equation (BE) [40]. While this is in principle better suited to model inhomogeneous systems, these earlier works did not provide a full account of finite size effects on mesoscopic length scales, nor did they establish the exact connection to the Kubo formula at a formal level.
@@
In this Letter, we provide a quantum kinetic theory of the linear response to a nearly constant and uniform electric field ${\bm{E}}$, for a multiband electron system, in presence of weak scalar disorder. With close connections with similar derivations in high-energy physics [41, & references therein], this is obtained from a controlled semiclassical expansion of the Schwinger-Keldysh theory [42, 43, 44], combined with a gauge invariant quantum geometric decomposition [45]. With the details the derivation to be found here [46], we expose our theory in terms of intraband and interband responses, uncovering mesoscopic gradients of the former as a previously overlooked source of the latter, which opens new areas of research. In the homogeneous infinite case, using Matsubara formalism [15, 16] , we establish the equivalence of our results with the Kubo formula with vertex correction in the ladder approximation. As a pedagogical example, we derive a theory of the spin-orbit torque (SOT) in the two-dimensional electron gas (2DEG), and uncover a new instance of intrinsic response as a Fermi surface property, reminiscent of the classical result obtained by Haldane in the case of the AHE [47]. Finally, we outline some implications for the theory of the recently observed orbital edge accumulation [48, 49] and orbital torque [50, & references therein], and we conclude by mentioning some future extensions.
@@
General formalism – Our main new result is a general expression for the interband part of the density matrix i.e., $\hat{\rho}_{inter}^{\scriptscriptstyle(1)}=\sum_{n\neq m}\hat{\rho}_{nm}^{\scriptscriptstyle(1)}$, in linear response to ${\bm{E}}$, with
$\!\!\!\hat{\rho}_{nm}^{\scriptscriptstyle(1)}$ $=-e\hbar\frac{f_{n}^{\scriptscriptstyle(eq)}-f_{m}^{\scriptscriptstyle(eq)}}{\varepsilon_{n}-\varepsilon_{m}}{\bm{E}}\cdot\hat{\mathcal{\bm{A}}}_{nm}-\hbar\partial_{\bm{x}}\frac{f_{n}^{\scriptscriptstyle(1)}+f_{m}^{\scriptscriptstyle(1)}}{2}\cdot\hat{\mathcal{\bm{A}}}_{nm}$
@@
×$\!\!\!\!\!\!\!\!\!\!\!\cdots\times\left[\delta\left(\underaccent{\tilde}{\varepsilon}_{q}\!-\!\varepsilon_{n}\right)\left(\underaccent{\tilde}{f}_{q}^{\scriptscriptstyle(1)}\!\!-\!\!f_{n}^{\scriptscriptstyle(1)}\right)+\delta\left(\underaccent{\tilde}{\varepsilon}_{q}\!-\!\varepsilon_{m}\right)\left(\underaccent{\tilde}{f}_{q}^{\scriptscriptstyle(1)}\!\!-\!\!f_{m}^{\scriptscriptstyle(1)}\right)\right],$
while the intraband part reads $\hat{\rho}_{intra}^{\scriptscriptstyle(1)}=\sum_{n}f^{\scriptscriptstyle(1)}_{n}\hat{P}_{n}$  [^1] This form of $\hat{\rho}^{\scriptscriptstyle(1)}$ is strictly valid for non degenerate bands. In the case of extended degeneracy, which is most generically realized in the form of Kramer’s degeneracy in time reversal and inversion symmetric systems, it remains valid if spin-orbit interaction can be safely neglected. Otherwise, a traceless part shall also be considered. In addition, in Eq.(1) we omit principal value terms which are usually interpreted as disorder-induced renormalizations of the Hamiltonian. Both can be found in [46]. and is controlled by a system of linear BEs
$\partial_{t}f_{n}^{\scriptscriptstyle(1)}+\partial_{\bm{p}}\varepsilon_{n}\cdot\partial_{\bm{x}}f_{n}^{\scriptscriptstyle(1)}+(e{\bm{E}})\cdot\partial_{\bm{p}}\varepsilon_{n}\partial_{\varepsilon}f_{n}^{\scriptscriptstyle(eq)}=C_{n}[{\bm{f}}^{\scriptscriptstyle(1)}].$ (2)
@@
To be more precise, $\hat{\rho}^{\scriptscriptstyle(1)}({\bm{x}},t;{\bm{p}})\equiv\hat{\rho}^{\scriptscriptstyle(1)}=\hat{\rho}^{\scriptscriptstyle(1)}_{intra}+\hat{\rho}^{\scriptscriptstyle(1)}_{inter}$ as determined by Eqs.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2) is the change in the gauge invariant Wigner function i.e., in the gauge invariant Wigner transform of the density matrix [52]. Once $\hat{\rho}^{\scriptscriptstyle(1)}$ is obtained, the linear response for any single particle observable $\smash{{\hat{O}}}$, in terms of its local density variation $\smash{o^{\scriptscriptstyle(1)}({\bm{x}},t)}$, is obtained from its Weyl symbol $\smash{{\hat{o}}({\bm{x}},{\bm{p}})}$ as $o^{\scriptscriptstyle(1)}({\bm{x}},t)=\int\frac{d^{d}p}{(2\pi\hbar)^{d}}\ \mathrm{Tr}[{\hat{\rho}}^{\scriptscriptstyle(1)}({\bm{x}},t;{\bm{p}})\ {\hat{o}}({\bm{x}},{\bm{p}})]$. Equations (Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2) determine the linear response of a non interacting system of electrons with charge $e(<0)$, specified by a multiband low energy Hamiltonian matrix i.e., ${\hat{h}}={\hat{h}}({\bm{p}})$, with ${\bm{p}}$ the kinetic crystal pseudo momentum. This matrix is acting on the Hilbert space of dimension $N$ of the internal degrees of freedom (spin, orbital), and it may equally correspond to a “toy” model or to a sophisticated Wannier tight-binding Hamiltonian [53, 54, 55]. It has in general $M\leq N$ distinct eigenvalues, the band energies, introduced as $\varepsilon_{n}\equiv\varepsilon_{n}({\bm{p}})$. The projectors into the corresponding eigenspaces, introduced as $\hat{P}_{n}\equiv\hat{P}_{n}({\bm{p}})$, read [56] $\hat{P}_{n}=\prod_{m}\frac{\hat{h}-\varepsilon_{m}\hat{\gamma}_{0}}{\varepsilon_{n}-\varepsilon_{m}}$, with $\hat{\gamma}_{0}$ the identity matrix. We further define $\hat{\mathcal{\bm{A}}}_{nm}={\bf i}\hat{P}_{n}\partial_{\bm{p}}\hat{P}_{m}$, which appears in the first line of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). In the special case of two non degenerate bands, it is immediate to verify that we have $(\hat{\mathcal{\bm{A}}}_{nm})^{i}={\mathcal{A}}^{i}_{nm}|n\rangle\langle m|$, with ${\mathcal{A}}^{i}_{nm}={\bf i}\langle n|\partial_{p^{i}}m\rangle$ a previously introduced abelian interband Berry connection [38, 57]. Hence, we recognize $\hat{\mathcal{\bm{A}}}_{nm}$ as a generalized non abelian interband Berry connection, defined in a manifestly $SU(N)$ gauge invariant way. Also appearing in Eqs.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems-2), are the linear changes in the intraband distribution functions i.e., $f_{n}^{\scriptscriptstyle(1)}\equiv f_{n}^{\scriptscriptstyle(1)}({\bm{x}},t;{\bm{p}})$, reducing in equilibrium to the Fermi-Dirac distribution function $f_{n}^{\scriptscriptstyle(eq)}\equiv n_{\scriptscriptstyle FD}(\varepsilon_{n})$. In the present work, we purposely limit our consideration of disorder to the most simple and generic model i.e., a dilute random distribution of scalar point scattering potentials treated at the level of the self-consistent second Born approximation after configuration averaging. It is parameterized by the density $n_{i}$ of defects and by a momentum independent scattering amplitude $v_{0}$. The collision integral appearing on the right-hand-side of the BE (2) reads
@@
$\!\!\!\!C_{n}[{\bm{f}}^{\scriptscriptstyle(1)}]\!=\!\!\sum_{q}\!\!\int\!\!\!\frac{d^{d}\underaccent{\tilde}{p}}{(2\pi\hbar)^{d}}\delta\left(\underaccent{\tilde}{\varepsilon}_{q}-\varepsilon_{n}\right)W_{nq}\left(\underaccent{\tilde}{f}_{q}^{\scriptscriptstyle(1)}-f_{n}^{\scriptscriptstyle(1)}\right)\!,$ (3)
@@
$W_{nq}\equiv W_{nq}({\bm{p}},\underaccent{\tilde}{\bm p})=\frac{2\pi n_{i}v_{0}^{2}}{\hbar}{\rm Tr}\left[\hat{P}_{n}\underaccent{\tilde}{\hat{P}}_{q}\right].$ (4)
@@
It has to be noted that we systematically use a tilde under-accent to indicate that a certain quantity depends on the integration momentum variable $\underaccent{\tilde}{\bm p}$. After introducing $\tau\propto\hbar/(n_{i}v_{0}^{2})$ as the relaxation time scale, we see that $\xi=1/\tau$ defines a natural scaling parameter for the disorder strength, while $\hbar$ is of course the formal expansion parameter for the quantum corrections to the classical kinetics [41, 45, & references therein].
@@
While the commonly considered decomposition of the Kubo-Streda type formulas into Fermi sea and Fermi surface terms [5, 7] is not physically significant in itself [47], the quantum kinetic approach leads naturally to a unique and gauge invariant split at the Wigner function level i.e., $\hat{\rho}^{\scriptscriptstyle(1)}=\hat{\rho}^{\scriptscriptstyle(1)}_{intra}+\hat{\rho}^{\scriptscriptstyle(1)}_{inter}$. Its physical meaning is further illuminated by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), establishing the interband response components as slave variables locally determined by the quantum geometry and the solutions of the BE (2). Hence, the quantum kinetic formulation establishes that interband contributions to observable densities are not quantities susceptible to advective or diffusive transport in real space. Considering the different terms on the right-hand side (RHS) of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), we recognize the first term as of purely quantum geometric origin and of order $\hbar\xi^{0}$, which will yield for instance the intrinsic contribution to the AHE [58, 11] which can be expressed in terms of the Berry curvature. If we look now at the third term on the RHS of our Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems), since $f_{n}^{\scriptscriptstyle(1)}\sim 1/\xi$, as it can be immediately verified by direct inspection of Eqs.(2-4), we see that it is also of order $\hbar\xi^{0}$ due to the compensating factor $(n_{i}v_{0}^{2})/\hbar\propto\xi$ in front of the momentum integral. Hence, even if this contribution is disorder mediated, it is ultimately independent of its strength. As demonstrated in the next section, this corresponds to the vertex correction in the ladder approximation to the interband part of the Kubo formula, for which the quantum kinetic approach provides an explicit expression at the Wigner function level, in terms of the solutions to the BE (2). As previously mentioned, even recent theoretical studies rely most often on a simplified version of the Kubo-Streda formula at the CBA level, with the leading disorder correction to the interband response found to be of order $\hbar\xi$ [59], which is qualitatively wrong and totally misses the true leading effect of disorder. The second term on the RHS of Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems) indicates that mesoscopic gradients in spatially inhomogeneous solutions of the BE (2) are also a source of interband quantum coherences, for which we provide here an explicit expression. To the best of our knowledge, this is an entirely new finding. Such gradients typically appear near boundaries and interfaces on the length scale of the electron mean-free-path $l_{e}\propto 1/\xi$ [60, 61, 62]. This opens a new area of research, since one can envision to engineer such gradients using nanofabrication techniques. Since $\hbar\partial_{\bm{x}}f_{n}^{\scriptscriptstyle(1)}\sim(\hbar/l_{e})f_{n}^{\scriptscriptstyle(1)}\sim\hbar\xi^{0}$, this is of the same order as the other contributions. Regarding the intraband response, it shall also be noted that there is no Berry curvature corrections to the BE (2), at the leading order $f_{n}^{\scriptscriptstyle(1)}\sim 1/\xi$. There are however sub-leading corrections at order $\hbar\xi^{0}$ which involve the quantum geometry, and are related to the intrinsic side-jump and skew-scattering contributions found with the semiclassical approach [35, 36, 37]. This is beyond the scope of the present letter, since a complete account requires to go beyond the ladder approximation [63], and will be provided within our general framework in a future work.
@@
Correspondence with the Kubo formula– It is well established that the intraband part of the linear response controlled by the BE (2), in the constant and homogeneous case, is identical to the intraband part of the Kubo formula with the vertex correction in the ladder approximation [15, 64]. In this section, we establish a similar equivalence for the interband response as given by Eq.(Quantum Kinetic Theory of the Linear Response for Weakly Disordered Multiband Systems). According to the Kubo formula at finite temperature in the constant and translationally invariant limits [15, 16], the expectation value of an observable ${\hat{o}}(\bm{p})\equiv{\hat{o}}$ reads $o^{\scriptscriptstyle(1)}=-eE^{i}\int_{p}\lim_{\omega\rightarrow 0}R^{i}(\omega)/({\bf i}\omega)$, with $\int_{p}\equiv\int\frac{{\rm d}^{d}{{p}}}{(2\pi\hbar)^{d}}$ and with $R^{i}({\bm{p}},\omega)\equiv R^{i}(\omega)$ being the analytical continuation (AC) of the Matsubara response function i.e., $R^{i}(\omega)={\cal R}^{i}(\omega_{\nu}\rightarrow-{\bf i}(\hbar\omega+{\bf i}0^{+}))$, with
@@
with $^{\hat{j}}^{i}=\partial_{p^{i}}\hat{h}$ the bare current vertex. From the corresponding expression of $o^{\scriptscriptstyle(1)}$ in terms of the density matrix, we immediately establish the correspondance
$\!\!{\hat{\rho}}^{\scriptscriptstyle(1)}_{nm}=-eE^{i}\!\lim_{\omega\rightarrow 0}\frac{\left[\sum_{\epsilon_{\mu}}\!\!\hat{P}_{n}\hat{\mathcal{G}}(\epsilon_{\mu}+\omega_{\nu}){\hat{J}}^{i}\hat{\mathcal{G}}(\epsilon_{\mu})\hat{P}_{m}\right]_{\scriptscriptstyle AC}}{{\bf i}\omega},$ (7)
@@
We can also write ${\hat{J}}^{i}=\sum_{n}J^{i}_{n}\hat{P}_{n}+\sum_{n\neq m}\hat{P}_{n}{\hat{J}}^{i}_{nm}\hat{P}_{m}$. At leading order in $\xi$, we then obtain the band projected part of Eq.(6) as